OpenStudy (anonymous):

dN/dt=r(1-N/K) solve for N(t) using Partial Fractions.

OpenStudy (anonymous):

\[\frac{dN}{dt}=r(1-\frac{N}{K})=r-\frac{rN}{K}\] If that's your problem, then it's a differential equation. I feel like that's not what you are trying to solve. Could you clarify if the quest is as such?

OpenStudy (anonymous):

how can i use partial fractions in this case to solve for N(t)

OpenStudy (anonymous):

This web sit does a good job of explaining what partial fractions is. http://www4.ncsu.edu/unity/lockers/users/f/felder/public/kenny/papers/partial.html You'll find out that this equation doesn't have a partial fraction.

OpenStudy (anonymous):

its an Allee Effect problem if you know how to do that one

OpenStudy (anonymous):

Yes, you solve it using a first order linear differntial equation. Have you taken differntial equation yet?

OpenStudy (anonymous):

not yet, need to do it for a biology question

OpenStudy (anonymous):

i know what the individual parts represent like what K is , what r is

OpenStudy (anonymous):

how do i solve it

OpenStudy (anonymous):

\[\frac{dN}{dt}=r(1-\frac{N}{K})=r-\frac{rN}{K}\] \[\frac{dN}{dt}+\frac{r}{K}N=r\] \[\mu*\frac{dN}{dt}+N\frac{d\mu}{dt}=\mu*r\] where \[\frac{r}{K}=\frac{1}{\mu}\frac{d\mu}{dt}\] Solve for that using separation of variables and integrating, you get. \[\mu=e^{\frac{r}{K}t}\] Essentially, \[\mu*N(t) = Ke^{\frac{r}{K}t}+c\] therefore \[N(t)=K+c*e^{\frac{r}{K}t}\] where c is a constant. If you don't follow, then I don't think this problem was properly assigned to you. Or this problem has more of a theoretical understanding rather than mathematical.

OpenStudy (anonymous):

If you look at the equation at the end, you'll see that it's a exponential rate. The population of N(t) grows at a rate of r/K as time increases.

OpenStudy (anonymous):

so if K=1 and r=1 and at N(0)=0.01 i can find c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get C=.99

OpenStudy (anonymous):

Did they really ask you to solve that problem in biology? That seems very farfetched. You must either have a hardcore teacher or misunderstood the problem. You C has the wrong sign. .01-1=-0.99

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

when i graphed my problem from 0 to 10 for t i got a negative exponetial graph

OpenStudy (anonymous):

yes because you are subtracting the exponent from your K value. Your initial condition made c negative. If your initial problem was dN/dt=r+rN/K (notice its positive instead of negative) then your solution would be N(t) = K + .99e^(rt/K)

OpenStudy (anonymous):

i tried to also find the equilibrium dN/dt=0. i got 2 but the question says 3 equilibrium points

OpenStudy (anonymous):

when dN/dt = 0 you can just solve the initial equation dN/dt = r - rN/K =0 where N=K

OpenStudy (anonymous):

I found the book where this guy got the question from its rediculous. he used a question from a diff eq book